# Calibrating Deep Neural Network using Euclidean Distance

### v1

@@ 7. Conclusions
In conclusion, our FCL effectively addresses the limitations of traditional Focal Loss by integrating a calibration term that significantly enhances the accuracy of probability estimations. Traditional Focal Loss has been instrumental in reducing misclassification rates by focusing on hard-to-classify samples; however, it often falls short in providing well-calibrated predictions. By ensuring that the loss function is strictly proper, FCL delivers more reliable class-posterior probabilities, which directly contributes to improved model confidence and reliability across various applications even in unforeseen circumstances.
@@ A.1. Evaluation Metrics
(18) $\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$
@@
$\text{acc}(B_{m})&=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})\hfill,\quad\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}\end{aligned}$
(20) $\text{MCE}=\max_{m\in{1,\ldots,M}}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|\end{aligned}\hfill$
(21) $\text{AdaECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|\end{aligned}\hfill$
$\text{ClasswiseECE}=\frac{1}{K}\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{|B_{km}|}{N_{k}}\left|\text{acc}(B_{km})-\text{conf}(B_{km})\right|\end{aligned}$
@@
A.2. Proof of Thm. 1 Focal Calibration Loss mitigates overconfidence (OC) and underconfidence (UC)
@@
We aim to bound the difference between the maximum predicted probability and the maximum true class probability. For any $\bm{x}\in\mathcal{X}$, we have:
(23) $\left|\max_{k}\hat{p}_{k}(\bm{x})-\max_{k}\eta_{k}(\bm{x})\right|\leq\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{\infty},$
@@
Let $k^{*}=\arg\max_{k}\hat{p}_{k}(\bm{x})$ and $k^{\dagger}=\arg\max_{k}\eta_{k}(\bm{x})$. Then,
(24) $\left|\max_{k}\hat{p}_{k}(\bm{x})-\max_{k}\eta_{k}(\bm{x})\right|=\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{\dagger}}(\bm{x})\right|.$
@@
(25) $\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{\dagger}}(\bm{x})\right|\leq\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{*}}(\bm{x})\right|+\left|\eta_{k^{*}}(\bm{x})-\eta_{k^{\dagger}}(\bm{x})\right|.$
Since $\eta_{k^{\dagger}}(\bm{x})\geq\eta_{k^{*}}(\bm{x})$ by the definition of $k^{\dagger}$ and $k^{*}$, it follows that:
(26) $\left|\eta_{k^{*}}(\bm{x})-\eta_{k^{\dagger}}(\bm{x})\right|=\eta_{k^{\dagger}}(\bm{x})-\eta_{k^{*}}(\bm{x})\geq 0.$
Similarly, since $\hat{p}_{k^{*}}(\bm{x})\geq\hat{p}_{k^{\dagger}}(\bm{x})$, we have:
(27) $\hat{p}_{k^{*}}(\bm{x})-\hat{p}_{k^{\dagger}}(\bm{x})\geq 0.$
@@
(28) $\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{\dagger}}(\bm{x})\right|\leq\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{*}}(\bm{x})\right|+\left(\eta_{k^{\dagger}}(\bm{x})-\eta_{k^{*}}(\bm{x})\right).$
Since $\eta_{k^{\dagger}}(\bm{x})-\eta_{k^{*}}(\bm{x})\leq\max_{k}\left|\eta_{k}(\bm{x})-\eta_{k}(\bm{x})\right|=0$, the second term is zero. Therefore,
(29) $\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{\dagger}}(\bm{x})\right|\leq\left|\hat{p}_{k^{*}}(\bm{x})-\eta_{k^{*}}(\bm{x})\right|.$
@@
(30) $\left|\max_{k}\hat{p}_{k}(\bm{x})-\max_{k}\eta_{k}(\bm{x})\right|\leq\max_{k}\left|\hat{p}_{k}(\bm{x})-\eta_{k}(\bm{x})\right|=\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{\infty}.$
@@
(31) $\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{\infty}\leq\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}.$
@@
(32) $\left|\max_{k}\hat{p}_{k}(\bm{x})-\max_{k}\eta_{k}(\bm{x})\right|\leq\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}.$
@@
(33) $\mathcal{L}_{\text{calib}}(f(\bm{x}),\bm{y})=\left\|\hat{\bm{p}}(\bm{x})-\bm{y}\right\|_{2}^{2}.$
Since $\bm{y}$ is a one-hot encoding sampled from the distribution $\bm{\eta}(\bm{x})$, we have:
(34) $\mathbb{E}_{\bm{y}\sim\bm{\eta}(\bm{x})}\left[\left\|\hat{\bm{p}}(\bm{x})-\bm{y}\right\|_{2}^{2}\right]=\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}^{2}+\sum_{k=1}^{K}\eta_{k}(\bm{x})(1-\eta_{k}(\bm{x})).$
The term $\sum_{k=1}^{K}\eta_{k}(\bm{x})(1-\eta_{k}(\bm{x}))$ is independent of $\hat{\bm{p}}(\bm{x})$ and represents the intrinsic uncertainty of the true distribution. Therefore, minimizing $\mathcal{L}_{\text{calib}}$ effectively minimizes $\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}^{2}$.
@@
By minimizing $\left\|\hat{\bm{p}}(\bm{x})-\bm{\eta}(\bm{x})\right\|_{2}^{2}$, we directly reduce the bound in Equation (32) on $\left|\max_{k}\hat{p}_{k}(\bm{x})-\max_{k}\eta_{k}(\bm{x})\right|$. This means that the difference between the highest predicted probability and the highest true probability decreases, mitigating both overconfidence and underconfidence.
Minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ includes minimizing the calibration loss component $\mathcal{L}_{\text{calib}}$, which reduces the Euclidean distance between the predicted probabilities $\hat{\bm{p}}(\bm{x})$ and the true class-posterior probabilities $\bm{\eta}(\bm{x})$. This, in turn, reduces the bound on the difference between the maximum predicted probability and the maximum true probability, thereby mitigating overconfidence and underconfidence.
A.3. Proof that Minimizing the Focal Calibration Loss Recovers the True Class-Posterior Probabilities
@@
Let $\mathcal{X}$ be the input space and $\mathcal{Y}=\{1,2,\dots,K\}$ the set of $K$ classes. For a given input $\bm{x}\in\mathcal{X}$, let $\bm{\eta}=[\eta_{1},\eta_{2},\dots,\eta_{K}]^{\top}$ denote the true class-posterior probabilities, where $\eta_{i}=\mathbb{P}(y=i\mid\bm{x})$ and $\sum_{i=1}^{K}\eta_{i}=1$; $\bm{q}=[q_{1},q_{2},\dots,q_{K}]^{\top}$ be the predicted probability vector, where $q_{i}\in(0,1)$ and $\sum_{i=1}^{K}q_{i}=1$. We aim to minimize the Focal Calibration Loss (FCL) and Our goal is to show that the minimizer $\bm{q}^{*}$ of $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ satisfies $q_{i}^{*}=\eta_{i}$ for all $i$.
(35) $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})=-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2},$
@@
Before proceeding, we establish that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ is strictly convex in $\bm{q}$ over the domain $q_{i}\in(0,1)$.
  • Calibration Loss Term: The term $\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}$ is strictly convex in $\bm{q}$ because it is a sum of squared terms.
  • Focal Loss Term: The term $-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}$ is convex in $q_{i}$ for $q_{i}\in(0,1)$. The function $\phi(q_{i})=-(1-q_{i})^{\gamma}\log q_{i}$ is convex in $q_{i}\in(0,1)$ for $\gamma\geq 0$. Since $\eta_{i}\geq 0$, and a non-negative weighted sum of convex functions is convex, the focal loss term is convex.
The sum of a strictly convex function (calibration loss) and a convex function (focal loss) is strictly convex (See Thm. 1 in Appendix A.4). Therefore, $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ is strictly convex in $\bm{q}$.
@@
Because $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ is strictly convex and the feasible set defined by $q_{i}>0$ and $\sum_{i=1}^{K}q_{i}=1$ is convex and closed, there exists a unique global minimizer $\bm{q}^{*}$.
@@
(36) $\mathcal{L}(\bm{q},\mu)=-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}+\mu\left(\sum_{i=1}^{K}q_{i}-1\right).$
We do not need to include inequality constraints $q_{i}>0$ explicitly because $q_{i}\in(0,1)$ due to the logarithm and the domain of $\bm{q}$.
@@
(37) $\frac{\partial\mathcal{L}}{\partial q_{i}}=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i})+\mu=0.$
@@
(38) $S_{i}(q_{i})=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i}).$
Then the stationarity condition becomes:$S_{i}(q_{i})+\mu=0.$ Since $\mu$ is the same for all $i$, we have:
(39) $S_{i}(q_{i})=S_{j}(q_{j})\quad\forall i,j.$
@@
(40) $S_{i}(\eta_{i})$ $=\eta_{i}\left[\gamma(1-\eta_{i})^{\gamma-1}\log\eta_{i}+\frac{(1-\eta_{i})^{\gamma}}{\eta_{i}}\right]+2\lambda(\eta_{i}-\eta_{i})$
(41) $=\eta_{i}\left[\gamma(1-\eta_{i})^{\gamma-1}\log\eta_{i}+\frac{(1-\eta_{i})^{\gamma}}{\eta_{i}}\right].$
@@
Since $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ is strictly convex, any stationary point is the unique global minimizer. Let’s evaluate $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ at $\bm{q}=\bm{\eta}$:
1. Calibration Loss term is minimized to 0: $\lambda\sum_{i=1}^{K}(\eta_{i}-\eta_{i})^{2}=0.$ 2. Focal Loss term $-\sum_{i=1}^{K}\eta_{i}(1-\eta_{i})^{\gamma}\log\eta_{i}$ is minimized when $q_{i}=\eta_{i}$ because the negative log-likelihood is minimized when the predicted probabilities match the true probabilities.
Since $\bm{q}=\bm{\eta}$ yields the minimal possible value of $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$, and the loss function is strictly convex, $\bm{q}^{*}=\bm{\eta}$ is the unique global minimizer.
@@
1. If $\eta_{i}=0$: The focal loss term for class $i$ becomes zero:$-\eta_{i}(1-q_{i})^{\gamma}\log q_{i}=0.$ The calibration loss term is $\lambda(q_{i}-0)^{2}=\lambda q_{i}^{2}$, minimized when $q_{i}=0$.
2. If $\eta_{i}=1$: The focal loss term becomes:$-(1-q_{i})^{\gamma}\log q_{i}.$ This is minimized when $q_{i}=1$. The calibration loss term $\lambda(q_{i}-1)^{2}$ is minimized when $q_{i}=1$.
@@
Therefore, minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\bm{q})$ over the probability simplex recovers the true class-posterior probabilities $\bm{q}^{*}=\bm{\eta}$.
A.4. Convexity of the Sum of Strictly Convex and Convex Functions
Theorem 1.

### v2

@@ 7 Conclusions
In conclusion, traditional Focal Loss has been instrumental in reducing misclassification rates by focusing on hard-to-classify samples; however, it often falls short in providing well-calibrated predictions. By ensuring that the loss function is strictly proper, FCL delivers more reliable class-posterior probabilities, which directly contributes to improved model confidence and reliability across various applications even in unforeseen circumstances.
8 Impact Statement
Ensuring accurate and trustworthy confidence scores in deep neural networks is vital for deploying them in high-stakes domains such as healthcare, autonomous driving, and finance. By introducing a focal calibration loss that leverages an explicit Euclidean distance term, our approach not only addresses class imbalance but also reduces miscalibration, aligning model outputs with real-world outcome frequencies. This dual improvement in accuracy and reliability fosters greater transparency and confidence when models are used to assist critical decisions. Consequently, our method’s blend of focal weighting and calibration stands to significantly enhance model trustworthiness across a range of practical applications.
@@ A.1 Evaluation Metrics
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (17)
@@
$\text{acc}(B_{m})$ $=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\mathbf{1}(\hat{y}_{i}=y_{i})\hfill,\quad\text{conf}(B_{m})=\frac{1}{|B_{m}|}\sum_{i\in B_{m}}\hat{p}_{i}$ (18)
$\begin{aligned} \text{MCE}=\max_{m\in{1,\ldots,M}}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|\end{aligned}\hfill$ (19)
$\begin{aligned} \text{AdaECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|\end{aligned}\hfill$ (20)
$\begin{aligned} \text{ClasswiseECE}=\frac{1}{K}\sum_{k=1}^{K}\sum_{m=1}^{M}\frac{|B_{km}|}{N_{k}}\left|\text{acc}(B_{km})-\text{conf}(B_{km})\right|\end{aligned}$ (21)
@@
A.2 Proof of Thm. 4.1 Focal Calibration Loss mitigates overconfidence (OC) and underconfidence (UC)
@@
We aim to bound the difference between the maximum predicted probability and the maximum true class probability. For any $\boldsymbol{x}\in\mathcal{X}$, we have:
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|\leq\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{\infty},$ (22)
@@
Let $k^{*}=\arg\max_{k}\hat{p}_{k}(\boldsymbol{x})$ and $k^{\dagger}=\arg\max_{k}\eta_{k}(\boldsymbol{x})$. Then,
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|=\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|.$ (23)
@@
$\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|\leq\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right|+\left|\eta_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|.$ (24)
Since $\eta_{k^{\dagger}}(\boldsymbol{x})\geq\eta_{k^{*}}(\boldsymbol{x})$ by the definition of $k^{\dagger}$ and $k^{*}$, it follows that:
$\left|\eta_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|=\eta_{k^{\dagger}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\geq 0.$ (25)
Similarly, since $\hat{p}_{k^{*}}(\boldsymbol{x})\geq\hat{p}_{k^{\dagger}}(\boldsymbol{x})$, we have:
$\hat{p}_{k^{*}}(\boldsymbol{x})-\hat{p}_{k^{\dagger}}(\boldsymbol{x})\geq 0.$ (26)
@@
$\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|\leq\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right|+\left(\eta_{k^{\dagger}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right).$ (27)
Since $\eta_{k^{\dagger}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\leq\max_{k}\left|\eta_{k}(\boldsymbol{x})-\eta_{k}(\boldsymbol{x})\right|=0$, the second term is zero. Therefore,
$\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{\dagger}}(\boldsymbol{x})\right|\leq\left|\hat{p}_{k^{*}}(\boldsymbol{x})-\eta_{k^{*}}(\boldsymbol{x})\right|.$ (28)
@@
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|\leq\max_{k}\left|\hat{p}_{k}(\boldsymbol{x})-\eta_{k}(\boldsymbol{x})\right|=\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{\infty}.$ (29)
@@
$\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{\infty}\leq\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}.$ (30)
@@
$\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|\leq\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}.$ (31)
@@
$\mathcal{L}_{\text{calib}}(f(\boldsymbol{x}),\boldsymbol{y})=\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{y}\right\|_{2}^{2}.$ (32)
Since $\boldsymbol{y}$ is a one-hot encoding sampled from the distribution $\boldsymbol{\eta}(\boldsymbol{x})$, we have:
$\mathbb{E}_{\boldsymbol{y}\sim\boldsymbol{\eta}(\boldsymbol{x})}\left[\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{y}\right\|_{2}^{2}\right]=\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}^{2}+\sum_{k=1}^{K}\eta_{k}(\boldsymbol{x})(1-\eta_{k}(\boldsymbol{x})).$ (33)
The term $\sum_{k=1}^{K}\eta_{k}(\boldsymbol{x})(1-\eta_{k}(\boldsymbol{x}))$ is independent of $\hat{\boldsymbol{p}}(\boldsymbol{x})$ and represents the intrinsic uncertainty of the true distribution. Therefore, minimizing $\mathcal{L}_{\text{calib}}$ effectively minimizes $\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}^{2}$.
@@
By minimizing $\left\|\hat{\boldsymbol{p}}(\boldsymbol{x})-\boldsymbol{\eta}(\boldsymbol{x})\right\|_{2}^{2}$, we directly reduce the bound in Equation (31) on $\left|\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\right|$. This means that the difference between the highest predicted probability and the highest true probability decreases, mitigating both overconfidence and underconfidence.
Minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}$ includes minimizing the calibration loss component $\mathcal{L}_{\text{calib}}$, which reduces the Euclidean distance between the predicted probabilities $\hat{\boldsymbol{p}}(\boldsymbol{x})$ and the true class-posterior probabilities $\boldsymbol{\eta}(\boldsymbol{x})$. This, in turn, reduces the bound on the difference between the maximum predicted probability and the maximum true probability, thereby mitigating overconfidence and underconfidence.
A.3 Proof that Minimizing the Focal Calibration Loss Recovers the True Class-Posterior Probabilities
@@
Let $\mathcal{X}$ be the input space and $\mathcal{Y}=\{1,2,\dots,K\}$ the set of $K$ classes. For a given input $\boldsymbol{x}\in\mathcal{X}$, let $\boldsymbol{\eta}=[\eta_{1},\eta_{2},\dots,\eta_{K}]^{\top}$ denote the true class-posterior probabilities, where $\eta_{i}=\mathbb{P}(y=i\mid\boldsymbol{x})$ and $\sum_{i=1}^{K}\eta_{i}=1$; $\boldsymbol{q}=[q_{1},q_{2},\dots,q_{K}]^{\top}$ be the predicted probability vector, where $q_{i}\in(0,1)$ and $\sum_{i=1}^{K}q_{i}=1$. We aim to minimize the Focal Calibration Loss (FCL) and Our goal is to show that the minimizer $\boldsymbol{q}^{*}$ of $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ satisfies $q_{i}^{*}=\eta_{i}$ for all $i$.
$\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})=-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2},$ (34)
@@
Before proceeding, we establish that $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex in $\boldsymbol{q}$ over the domain $q_{i}\in(0,1)$.
  • Calibration Loss Term: The term $\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}$ is strictly convex in $\boldsymbol{q}$ because it is a sum of squared terms.
  • Focal Loss Term: The term $\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}$ is convex in $q_{i}$ for $q_{i}\in(0,1)$. The function $\phi(q_{i})=-(1-q_{i})^{\gamma}\log q_{i}$ is convex in $q_{i}\in(0,1)$ for $\gamma\geq 0$. Since $\eta_{i}\geq 0$, and a non-negative weighted sum of convex functions is convex, the focal loss term is convex.
The sum of a strictly convex function (calibration loss) and a convex function (focal loss) is strictly convex (See Thm. A.1 in Appendix A.4). Therefore, $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex in $\boldsymbol{q}$.
@@
Because $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex and the feasible set defined by $q_{i}>0$ and $\sum_{i=1}^{K}q_{i}=1$ is convex and closed, there exists a unique global minimizer $\boldsymbol{q}^{*}$.
@@
$\mathcal{L}(\boldsymbol{q},\mu)=-\sum_{i=1}^{K}\eta_{i}(1-q_{i})^{\gamma}\log q_{i}+\lambda\sum_{i=1}^{K}(q_{i}-\eta_{i})^{2}+\mu\left(\sum_{i=1}^{K}q_{i}-1\right).$ (35)
We do not need to include inequality constraints $q_{i}>0$ explicitly because $q_{i}\in(0,1)$ due to the logarithm and the domain of $\boldsymbol{q}$.
@@
$\frac{\partial\mathcal{L}}{\partial q_{i}}=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i})+\mu=0.$ (36)
@@
$S_{i}(q_{i})=\eta_{i}\left[\gamma(1-q_{i})^{\gamma-1}\log q_{i}+\frac{(1-q_{i})^{\gamma}}{q_{i}}\right]+2\lambda(q_{i}-\eta_{i}).$ (37)
[…]
$S_{i}(q_{i})=S_{j}(q_{j})\quad\forall i,j.$ (38)
@@
$S_{i}(\eta_{i})$ $=\eta_{i}\left[\gamma(1-\eta_{i})^{\gamma-1}\log\eta_{i}+\frac{(1-\eta_{i})^{\gamma}}{\eta_{i}}\right]+2\lambda(\eta_{i}-\eta_{i})$ (39)
$=\eta_{i}\left[\gamma(1-\eta_{i})^{\gamma-1}\log\eta_{i}+\frac{(1-\eta_{i})^{\gamma}}{\eta_{i}}\right].$ (40)
@@
Since $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ is strictly convex, any stationary point is the unique global minimizer. Let’s evaluate $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ at $\boldsymbol{q}=\boldsymbol{\eta}$:
1. Calibration Loss term is minimized to 0: $\lambda\sum_{i=1}^{K}(\eta_{i}-\eta_{i})^{2}=0.$ 2. Focal Loss term $\sum_{i=1}^{K}\eta_{i}(1-\eta_{i})^{\gamma}\log\eta_{i}$ is minimized when $q_{i}=\eta_{i}$ because the negative log-likelihood is minimized when the predicted probabilities match the true probabilities.
Since $\boldsymbol{q}=\boldsymbol{\eta}$ yields the minimal possible value of $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$, and the loss function is strictly convex, $\boldsymbol{q}^{*}=\boldsymbol{\eta}$ is the unique global minimizer.
@@
[…]
@@
Therefore, minimizing the Focal Calibration Loss $\mathcal{L}_{\text{FCL}}^{\gamma,\lambda}(\boldsymbol{q})$ over the probability simplex recovers the true class-posterior probabilities $\boldsymbol{q}^{*}=\boldsymbol{\eta}$.
A.4 Convexity of the Sum of Strictly Convex and Convex Functions
Theorem A.1.